Hillinger, Inc. issues $100,000 of 4% bonds on January 1, Year 1. The bonds have a five-year term and pay interest semiannually on June 30 and December 31 each year. Assuming a market interest rate of 5%, what is the carrying value of the bonds as of December 31, Year 1? A.$96,015 B.$96,415 C.$97,000
Answer #1
Cash interest Interest expense Discount amortized Carrying value
January 1, Year 1 95624
June 30, Year 1 2000 2391 391 96015
December 31, Year 1 2000 2400 400 96415
Option B $96,415 is correct
Workings:
Semi-annual interest 2000 =100000*4%*6/12
Number of periods 10 =5*2
Amount PV factor 2.5% Present value
Semi-annual interest 2000 8.75206 17504
Principal 100000 0.78120 78120
Total 95624
PV factor 2.5%
Semi-annual interest 8.75206 =(1-(1.025)^-10)/0.025
Principal 0.78120 =1/1.025^10
Interest expense
June 30, Year 1 2391 =95624*5%/2
December 31, Year 1 2400 =96015*5%/2